# LocDiffusion: Identifying Locations on Earth by Diffusing in the Hilbert Space

### v1

@@ Abstract
Image geolocalization is a fundamental yet challenging task, aiming at inferring the geolocation on Earth where an image is taken. Existing methods approach it either via grid-based classification or via image retrieval. Their performance significantly suffers when the spatial distribution of test images does not align with such choices. To address these limitations, we propose to leverage diffusion as a mechanism for image geolocalization. To avoid the problematic manifold reprojection step in diffusion, we developed a novel spherical positional encoding-decoding framework called Spherical Harmonics Dirac Delta (SHDD), which encodes points on a spherical surface (e.g., geolocations on Earth) into a Hilbert space of Spherical Harmonics coefficients and decodes points (geolocations) by mode-seeking. We also propose a novel SirenNet-based architecture called CS-UNet to learn the conditional backward process in the latent SHDD space by minimizing a latent KL-divergence loss. We train a conditional latent diffusion model called LocDiffusion that generates geolocations under the guidance of images – to the best of our knowledge, the first generative model for image geolocalization by diffusing geolocation information in a hidden location embedding space. We evaluate LocDiffusion against SOTA image geolocalization baselines. It achieves competitive geolocalization performance and demonstrates significantly stronger generalizability to unseen geolocations.
Machine Learning, ICML
@@ 1 Introduction
Predicting locations on Earth based on a given condition (e.g., input image or text) is a fundamental yet challenging task. Image geolocalization, being a prominent example of this task, aims at predicting the geolocations only based on images, such as wildlife photos, street views, and remote sensing images. However, unlike image classification, solutions to image geolocalization are less mature because its ground-truths are locations represented by real-valued coordinates on the spherical surface. While regression models are commonly used to predict real-valued labels, they are proved to be tricky to train and perform especially poorly on image geolocalization due to the highly complex and non-linear mapping between the image space and the geospatial space (Vo et al., 2017; Izbicki et al., 2020). As an alternative solution, researchers employ pre-defined geographical classes (e.g. divide Earth into disjoint or hierarchical grid cells) or geo-tagged image galleries (e.g. a set of reference geotagged images) to map the real-valued ground-truth coordinates to discrete labels (e.g. the ID of the grid cell the ground-truth falls into or the ID of the reference image in the gallery that has the closest geotag as the label of a ground-truth location), subsequently transforming image geolocalization problem into a special case of image classification or image-image/image-location retrieval task. For example, (Vo et al., 2017), CPlaNet (Seo et al., 2018), and PIGEON (Haas et al., 2024) partition the Earth’s surface into non-overlapping cells and convert the image geolocalization problem into an image classification problem. GeoCLIP (Vivanco et al., 2023) uses a contrastive learning framework to align pretrained image embeddings with geographical location embeddings in the gallery and achieves competitive performance on two widely used datasets. However, the spatial resolution of these approaches is constrained by the size of the cells or the spatial distribution of gallery images/locations.
Figure 1: Latent diffusion for image geolocalization. Black solid/dotted arrows denote encoding/decoding steps. Orange modules are learnable, while blue modules are deterministic with no learning parameters. (1) It is difficult to diffuse in the position encoding space (Mai et al., 2023b) because valid positional encodings are also sparse which leads to difficulties in diffusion model training and decoding. (2) The locational embedding space is dense and can perform diffusion processes, but the non-linear mapping between the position encoding and location embedding space makes decoding back to a correct coordinates extremely difficult. Minimizing distances in the location embedding space may not minimize geographic distance. (3) The SHDD encoding space is dense – every point $\mathbf{e}$ in this encoding space corresponds to a spherical function $F_{\mathbf{e}}$, whose difference from the spherical Dirac delta function $\delta_{(\theta_{0},\phi_{0})}$ of the ground truth location $(\theta_{0},\phi_{0})$ is measured by the reverse KL-divergence $\mathcal{E}$. (4) The SHDD decoding addresses the non-linearity problem. The heatmaps (4a), (4b) represents the distance from the spherical point represented by the embedding/encoding to the yellow star point in the middle. The distance measured by SHDD is significantly smoother.
@@
Diffusion models have demonstrated great capacity in directly generating continuous outputs such as images and modeling their complex distributions. They are commonly applied to points in Euclidean spaces (Song et al., 2020; Ho et al., 2020b; Song et al., 2021) or the geometric structures defined in Euclidean spaces (Xu et al., 2023). This motivates us to develop diffusion-based image geolocalization methods that output location predictions on the spherical surface with a finer spatial resolution by using multi-scale location representations (Mai et al., 2020a, 2023a; Rußwurm et al., 2024) and without dependence on predefined grid cells or galleries. However, naively performing diffusion in the coordinate space faces two major drawbacks. First, geographical locations do not form an Euclidean space. They reside on an embedded Riemannian manifold [^1] Geographical locations are distributed on a 2-dimensional Riemannian manifold (i.e., the sphere surface) embedded in the 3-dimensional Euclidean space.. Diffusion in the geographical coordinate space is ineffective because of projection distortion and sparsity, i.e., performing diffusion on the XYZ coordinates will likely lead to a point that is not on the spherical surface. It is possible to perform diffusion on the manifolds, but it is very computationally expensive (Huang et al., 2022). Second, more importantly, raw coordinates cannot represent rich multi-scale geographical information or modeling complex spatial distributions (Mai et al., 2023b; Rußwurm et al., 2024). Diffusion in the coordinate space would require non-standard diffusion model with multi-scale representations internally.
In order to achieve good modeling power for complex distributions over space, location representation methods (Mac Aodha et al., 2019; Mai et al., 2020b, 2022, 2023b; Rußwurm et al., 2024; Wu et al., 2024) commonly adopt a multi-scale position encoding with deterministic transformations followed by a learnable location embedding layers. However, despite their wide applicability, neither the position encoding nor the location embedding space is suitable for developing a location diffusion model due to sparsity problem during diffusion and non-linearity problem during decoding as illustrated in Figure 1(1)(2). On the one hand, the position encoding space has a “sparsity problem”. The position encoding layer commonly increase the dimensionality of representations significantly from the coordinate space. Therefore, all valid positional encodings form a low-dimensional manifold embedded in a high-dimensional Euclidean space. If we diffuse in the embedded high-dimensional Euclidean space and train a decoder to map Euclidean points back to geographic coordinates, the sparsity of position encodings (from the available training data) makes it difficult to learn smooth local interpolations that can generalize to unseen data. On the other hand, while location embedding space is dense and suitable for forward and backward diffusion processes, it is difficult to learn an inverse mapping which decodes the location embedding directly back to the coordinate space, skipping the position encoding space, because of the non-linear mapping between position encodings and location embeddings.
@@
We hypothesize that the ideal space to develop latent diffusion models for spherical location generation should be both dense and easy to find projections back to the coordinate space. Motivated by this observation, we propose a novel spherical position encoding method called Spherical Harmonics Dirac Delta (SHDD) Representation. Figure 1(3)(4) illustrates how our method addresses the sparsity problem by encoding a spherical point $(\theta_{0},\phi_{0})$ as a spherical Dirac delta function $\delta_{(\theta_{0},\phi_{0})}$. In the SHDD encoding space, every point $\mathbf{e}$ uniquely corresponds to a spherical function $F_{\mathbf{e}}$ and can be seen as an approximation of a spherical Dirac delta function. The level of noise $\mathcal{E}$ can be continuously measured by the reverse KL-divergence between $F_{\mathbf{e}}$ and $\delta_{(\theta_{0},\phi_{0})}$. Then the latent diffusion in the SHDD encoding space equals gradually adding noise to the ground-truth $\delta_{(\theta_{0},\phi_{0})}$ (forward process) and find a sequence of $F_{\mathbf{e}}$ that gradually reduce $\mathcal{E}$ (backward process). During decoding, the learning-free SHDD Decoder evaluates the corresponding spherical function $F_{\mathbf{e}}$ and decodes it as the spherical point whose corresponding spherical Dirac delta function minimizes $\mathcal{E}$. Figure 1 4(b) demonstrates that our SHDD encoding space shows less decoding non-linearity than existing location representation learning methods such as Sphere2Vec (Mai et al., 2023b) and (Rußwurm et al., 2024). Therefore, diffusion in SHDD encoding space will be more stable and easier to converge.
@@
Equipped with the Hilbert (i.e. infinite dimensional Euclidean) SHDD encoding space and the SHDD decoder, we can now perform conventional latent diffusion for location generation. We propose a novel SirenNet-based architecture called Conditional Siren-UNet (CS-UNet) to learn the conditional backward diffusion process, i.e, to generate spherical points from random Gaussian noise given conditions such as images and texts. We call the integrated framework, including SHDD encoding, CS-UNet latent diffusion, and SHDD decoding, which enables efficient conditional generation of spherical points, the LocDiffusion model. Compared to state-of-the-art models on global image geolocalization tasks, LocDiffusion is proven to be more spatially generalizable than existing retrieval-based geolocalization models by ablation experiments.
@@ 2 Related Work
Traditional geolocalization methods typically employ either a classification approach or an image retrieval approach. The former divides the Earth’s surface into non-overlapping or hierarchical cells and classifies images accordingly (Pramanick et al., 2022a; Vo et al., 2017; Muller-Budack et al., 2018; Haas et al., 2024) while the later approach identifies the location of a given image by matching it with a database of image-location pairs (Shi et al., 2020; Zhu et al., 2023; Zhou et al., 2024). Using fewer cells results in lower location prediction accuracy while using smaller cells reduces the number of training examples per class and risks overfitting (Seo et al., 2018). On the other hand, retrieval-based systems usually suffer from poor search quality and inadequate coverage of the global geographic landscape.
Diffusion in the coordinate space.
Conventional diffusion models cannot function well in the spherical coordinate space (e.g., 3D coordinates representing points on a sphere) because valid points for diffusion are too sparse, i.e., adding or removing noise to a point on the spherical manifold almost always results in a point outside the manifold. While certain coordinates such as latitude and longitude can remain in the valid manifold with noises, these spaces are non-Euclidean and not suitable for existing denoising diffusion models (Song et al., 2021). They can also cause significant distortions in localization (e.g., polar areas).
Riemannian diffusion models.
There are two common strategies to address the above problem. The first strategy is to project a point on the sphere to its tangent space (an Euclidean space), add/remove noise in the tangent space, and re-project the noised/denoised point in the tangent space back to the surface (Rozen et al., 2021). The second strategy is to derive formulas for direct Riemannian diffusion (Huang et al., 2022). The main drawback of both strategies is their computational complexity. In the first case, each projection operation takes time, making acceleration based on DDIM (Song et al., 2021) impossible, because the projections are accurate only when the diffusion steps are adequately small. For the second case, the Riemannian diffusion formulation is much more complicated than the Euclidean version. The model architectures, training tricks, and other useful techniques developed for conventional diffusion models can not be easily transferred.
@@
The distinction between positional encoding and location embedding lies in semantics: the positional encoding is only a task-agnostic transformation of the coordinates $\mathbf{x}$, but the location embedding carries task-specific information. For example, it can contain information about spatial distributions of species if trained on geo-aware species fine-grained recognition tasks (Mac Aodha et al., 2019; Mai et al., 2023b, a; Cole et al., 2023). Some prior work, e.g., NeRF (Mildenhall et al., 2020), utilized positional encoding to represent location information. This task-agnostic method focuses on capturing the position or order of elements within a sequence. In contrast, many location encoders are specifically designed to capture context-aware location information (Mai et al., 2020a). Please refer to Appendix A.1 and Mai et al. (2023b) for more details.
@@ 3 Preliminaries
3.1 Real Basis of Spherical Harmonics
Let $\mathbf{p}=(\theta,\phi)$ be a location on the spherical surface using angular coordinates where $\theta\in[0,\pi)$ and $\phi\in[0,2\pi)$. For any function $F(\theta,\phi)$ on the sphere, there exists a unique infinite-dimensional real-valued vector of coefficients $\{C_{lm}\}$ (we may call it coefficient vector) such that
$\forall(\theta,\phi),F(\theta,\phi)=\sum_{l=0}^{\infty}\sum_{m=-l}^{l}C_{lm}Y_{lm}(\theta,\phi)$ (1)
where $l$ is called degree and $m$ is called order and $Y_{lm}(\theta,\phi)$ is the real basis of spherical harmonics at degree $l$ and order $m$. The detailed computation of $Y_{lm}$ can be found in Appendix A.2. In this way, any function on the sphere can be uniquely represented by its coefficient vector.
3.2 Spherical Dirac Delta Function
Conventionally, a Dirac delta function $\delta$ is defined as a distribution on the real line where all probability mass concentrates on one single value, i.e., a single-point distribution. Analogously, a spherical Dirac delta function is a probability density function over the spherical surface whose mass all concentrates on one point:
$\delta_{(\theta_{0},\phi_{0})}(\theta,\phi)=\begin{cases}\infty&\theta=\theta_{0},\phi=\phi_{0}\\ 0&\text{otherwise}\end{cases}$ (2)
Therefore, we use a spherical Dirac delta function to uniquely represent any point $(\theta_{i},\phi_{i})$ on the sphere by mapping it to $\delta_{(\theta_{i},\phi_{i})}$. Representing a point as a function allows us to use spherical harmonics to represent points on the spherical surface.
4 LocDiffusion Framework
In this section, we will introduce the theory and techniques we employ in our LocDiffusion model that enable spherical location generation via latent diffusion. Our aim is to find a position encoding space that does not suffer from the sparsity problem and the non-linearity problem so we can efficiently perform latent diffusion. We first analyze what properties we need to achieve this and propose the Spherical Harmonics Dirac Delta (SHDD) Encoding-Decoding framework accordingly. Then we prove that SHDD satisfies all the desired properties. Following that, we propose the Conditional Siren-UNet (CS-UNet) architecture to learn the conditional backward process for latent diffusion. We also develop computational techniques based on the properties of SHDD representation so that the training and inference of LocDiffusion are efficient.

### v2

# LocDiff: Identifying Locations on Earth by Diffusing in the Hilbert Space

@@ Abstract
Image geolocalization is a fundamental yet challenging task, aiming at inferring the geolocation on Earth where an image is taken. State-of-the-art methods employ either grid-based classification or gallery-based image-location retrieval, whose spatial generalizability significantly suffers if the spatial distribution of test images does not align with the choices of grids and galleries. Recently emerging generative approaches, while getting rid of grids and galleries, use raw geographical coordinates and suffer quality losses due to their lack of multi-scale information. To address these limitations, we propose a multi-scale latent diffusion model called LocDiff for image geolocalization. We developed a novel positional encoding-decoding framework called Spherical Harmonics Dirac Delta (SHDD) Representations, which encodes points on a spherical surface (e.g., geolocations on Earth) into a Hilbert space of Spherical Harmonics coefficients and decodes points (geolocations) by mode-seeking on spherical probability distributions. We also propose a novel SirenNet-based architecture (CS-UNet) to learn an image-based conditional backward process in the latent SHDD space by minimizing a latent KL-divergence loss. To the best of our knowledge, LocDiff is the first image geolocalization model that performs latent diffusion in a multi-scale location encoding space and generates geolocations under the guidance of images. Experimental results show that LocDiff can outperform all state-of-the-art grid-based, retrieval-based, and diffusion-based baselines across 5 challenging global-scale image geolocalization datasets, and demonstrates significantly stronger generalizability to unseen geolocations.
Figure 1: Multi-scale latent diffusion for image geolocalization. Black solid/dotted arrows denote encoding/decoding steps. Orange modules are learnable, while blue modules are deterministic with no learning parameters. (1) It is difficult to diffuse in the position encoding space (mai2023sphere2vec, ) because valid positional encodings are sparse which leads to difficulties in diffusion model training and decoding. (2) The locational embedding space is dense and can perform diffusion processes, but the non-linear mapping between the position encoding and location embedding space makes decoding back to a correct coordinates extremely difficult. Minimizing distances in the location embedding space may not minimize geographic distance. (3) The SHDD encoding space is dense – every point $\mathbf{e}$ in this encoding space corresponds to a spherical function $F_{\mathbf{e}}$, whose difference from the spherical Dirac delta function $\delta_{(\theta_{0},\phi_{0})}$ of the ground truth location $(\theta_{0},\phi_{0})$ is measured by the reverse KL-divergence $\mathcal{E}$. (4) The SHDD decoding addresses the non-linearity problem. The heatmaps (4a), (4b) represents the distance from the spherical point represented by the embedding/encoding to the yellow star point in the middle. The distance measured by SHDD is significantly smoother.
@@ 1 Introduction
Location decoding, i.e., predicting geolocations from given context, remains a challenging and rarely studied problem which has potential applications in numerous tasks including trajectory synthesis rao2023cats , building footprint segmentation hu2023polybuilding , and the widely studied image geolocalization tasks seo2018cplanet ; geoclip . As one of the representative tasks for location decoding, image geolocalization aims at predicting locations on Earth based on a given image, which potentially ranges from many types such as wild life photos, street views, and remote sensing images. However, unlike image classification, solutions to image geolocalization are less mature because their ground-truths are locations represented by real-valued coordinates on the spherical surface. While regression models are commonly used to predict real-valued labels, they are proven to be tricky to train and perform especially poorly on the global-scale image geolocalization problem due to the highly complex and non-linear mapping between the image space and the geospatial space (vo2017revisiting, ; izbicki2020exploiting, ). As an alternative transductive solution, researchers employ pre-defined geographical classes (e.g. divide Earth into disjoint or hierarchical grid cells) (vo2017revisiting, ; seo2018cplanet, ) or geo-tagged image galleries (e.g. a set of reference geotagged images) (geoclip, ) to map the real-valued ground-truth coordinates to discrete labels (e.g. the ID of the grid cell or the ID of the reference image in the gallery), subsequently transforming the image geolocalization problem into a special case of image classification or image-image/image-location retrieval task. For example, [L]kNN vo2017revisiting , CPlaNet (seo2018cplanet, ), and PIGEON (Haas_2024_CVPR, ) partition the Earth’s surface into non-overlapping cells and convert the image geolocalization problem into an image classification problem. GeoCLIP (geoclip, ) uses a contrastive learning framework to align pretrained image embeddings with geographical location embeddings in the gallery. However, the spatial resolution of these approaches is constrained by the size of the cells or the availability of gallery images/locations.
@@
Generative models such as diffusion models ho2020denoising ; song2021denoising , score-based generative models song2020score , and flow matching models lipman2023flow , have demonstrated great capacity in directly generating continuous outputs such as images and modeling their complex distributions. They are commonly applied to points in Euclidean spaces (song2020score, ; ho2020denoising, ; song2021denoising, ) or the geometric structures defined in Euclidean spaces (xu2023geometric, ). This motivates us to develop diffusion-based image geolocalization methods that decode locations on the spherical surface with a finer spatial resolution by using multi-scale location representations(mai2020space2vec, ; mai2023csp, ; russwurm2024locationencoding, ) without dependence on predefined grid cells or galleries (vo2017revisiting, ; seo2018cplanet, ; geoclip, ; Haas_2024_CVPR, ). However, simply performing diffusion or flow matching in the original coordinate space faces two major challenges. First, geographical locations reside on an embedded Riemannian manifold instead of an Euclidean space [^1] Geographical locations are distributed on a 2-dimensional Riemannian manifold (i.e., the sphere surface) embedded in the 3-D Euclidean space.. This makes noise adding/removal problematic. If we use coordinates on the manifold such as longitudes and latitudes, the noise adding/removal will lead to projection distortions, especially near the polar areas. If we use coordinates in the 3-D Euclidean space, every forward/backward step will likely output points that are not on the spherical surface dufour2024world80timestepsgenerative , which requires noise adding/removal in the tangent space and re-projection back to the spherical surface, such as in Diff $\mathbb{R}^{3}$ dufour2024world80timestepsgenerative and FM $\mathbb{R}^{3}$ dufour2024world80timestepsgenerative . The re-projection operation introduces extra computational costs. Secondly, and more importantly, raw coordinates lack multi-scale information which is necessary for modeling complex spatial distributions (mai2023sphere2vec, ; russwurm2024locationencoding, ). This practice can yield suboptimal geolocalization results. As later shown in Table 2, compared with generative models operating directly on raw coordinates (e.g., Diff $\mathbb{R}^{3}$, FM $\mathbb{R}^{3}$ and RFM $\mathcal{S}_{2}$ dufour2024world80timestepsgenerative ), our multi-scale latent location diffusion model LocDiff demonstrates superior performance on 2 image geolocalization datasets across all evaluation metrics measured in various spatial scales.
In order to achieve multi-scale modeling power for complex distributions over space, neural location encoding methods (space2vec_iclr2020, ; mai2022review, ; mai2023sphere2vec, ; mai2023csp, ; klemmer2025satclip, ; russwurm2024locationencoding, ; wu2024torchspatial, ; yang2022dynamic, ) commonly adopt a multi-scale position encoding with deterministic transformations followed by a learnable location embedding layer. However, despite their wide applicability, neither the position encoding nor the location embedding space is suitable for developing a location diffusion model due to sparsity problem during diffusion and non-linearity problem during decoding as illustrated in Figure 1(1) and (2). Therefore, we hypothesize that
The ideal space to develop latent diffusion models for spherical location generation should be both dense and easy to find projections back to the coordinate space.
@@
Motivated by this, we propose a novel spherical position encoding method called Spherical Harmonics Dirac Delta (SHDD) Representation. Figure 1(3) and (4) illustrates how our method addresses the sparsity problem by encoding a spherical point $(\theta_{0},\phi_{0})$ as a spherical Dirac delta function $\delta_{(\theta_{0},\phi_{0})}$. In the SHDD encoding space, every point $\mathbf{e}$ uniquely corresponds to a spherical function $F_{\mathbf{e}}$ and can be seen as an approximation of a spherical Dirac delta function. The level of noise $\mathcal{E}$ can be continuously measured by the reverse KL-divergence between $F_{\mathbf{e}}$ and $\delta_{(\theta_{0},\phi_{0})}$. Then the latent diffusion in the SHDD encoding space equals gradually adding noise to the ground-truth $\delta_{(\theta_{0},\phi_{0})}$ (forward process) and finding a sequence of $F_{\mathbf{e}}$ that gradually reduces $\mathcal{E}$ (backward process). During decoding, the learning-free SHDD Decoder evaluates the corresponding spherical function $F_{\mathbf{e}}$ and decodes it as the spherical point whose corresponding spherical Dirac delta function minimizes $\mathcal{E}$. Figure 1 4(b) demonstrates that our SHDD encoding space shows less decoding non-linearity than existing location representation learning methods such as Sphere2Vec (mai2023sphere2vec, ) and SH russwurm2024locationencoding . Therefore, diffusion in the SHDD encoding space will be more stable and easier to converge.
@@
Equipped with the Hilbert (i.e. infinite-dimensional Euclidean) SHDD encoding space and the SHDD decoder, we can now perform conventional latent diffusion for location generation. We propose a novel SirenNet-based architecture called Conditional Siren-UNet (CS-UNet) to learn the conditional backward diffusion process, i.e, to generate spherical points from random Gaussian noise given conditions such as images and texts. We call the integrated framework, including SHDD encoding, CS-UNet latent diffusion, and SHDD decoding, the LocDiff model. Results show that LocDiff outperforms all state-of-the-art models on 5 global-scale image geolocalization datasets at large scales (250 km, 750 km, 2500 km), and by combining retrieval-based models (See Appendix A.6), the hybrid variant LocDiff-H achieves superior performance at small scales (1 km, 25 km). LocDiff also outperforms the recently proposed generative geolocalization model RFM $\mathcal{S}_{2}$ dufour2024world80timestepsgenerative , which shows the advantages of our latent diffusion approach over generation in the original coordinate space. Moreover, we demonstrate that LocDiff is more spatially generalizable to unseen locations than retrieval-based models such as GeoCLIP (geoclip, ).
@@ 2 Related Work
Traditional geolocalization methods typically employ either a classification approach or an image retrieval approach. The former divides the Earth’s surface into non-overlapping or hierarchical cells and classifies images accordingly (Pramanick22, ; vo2017revisiting, ; muller2018geolocation, ; Haas_2024_CVPR, ) while the latter approach identifies the location of a given image by matching it with a database of image-location pairs (shi2020looking, ; zhu2023r2former, ; zhou2024img2loc, ; geoclip, ). Using fewer cells results in lower location prediction accuracy while using smaller cells reduces the number of training examples per class and risks overfitting (seo2018cplanet, ). On the other hand, retrieval-based systems usually suffer from poor search quality and inadequate coverage of the global geographic landscape.
Geolocalization by diffusion and flow matching in the original coordinate space.
Dufour et al dufour2024world80timestepsgenerative recently proposed a set of generative image geolocalization models on the original coordinate space, including Diff $\mathbb{R}^{3}$, FM $\mathbb{R}^{3}$, and RFM $\mathcal{S}_{2}$. The first two models perform diffusion and flow matching on the 3D Euclidean space respectively and project the output back on the sphere. RFM $\mathcal{S}_{2}$, instead, performs flow matching on the spherical Riemannian manifold by projecting back and forth to the tangent space of each location. However, operating on the original coordinate space cannot capture the rich multi-scale geographic information required for modeling complex distributions on the spherical surface. Our LocDiff overcomes this limitation by performing diffusion in a latent space which captures rich multi-scale location information.
Riemannian generative models.
Conventional generative models cannot function well in the spherical coordinate space (e.g., 3D coordinates representing points on a sphere) due to the projection distortion and sparsity issue. There are two common strategies to address this problem. The first strategy is to project a point on the sphere to its tangent space (an Euclidean space), add/remove noise in the tangent space, and then spherical re-project the noised/denoised point in the tangent space back to the surface (rozen2021moser, ) like RFM $\mathcal{S}_{2}$ dufour2024world80timestepsgenerative does. The second strategy is to derive formulas for direct Riemannian diffusion (huang2022riemannian, ). The main drawback of both strategies is their computational complexity. In the first strategy, each projection operation takes time, making sampling acceleration based on DDIM (song2021denoising, ) impossible, because the projections are accurate only when the diffusion steps are adequately small. For the second strategy, the Riemannian diffusion formulation is much more complicated than the Euclidean version. The model architectures, training tricks, and other useful techniques developed for conventional diffusion models can not be easily transferred.
@@
The distinction between positional encoding and location embedding lies in semantics: the positional encoding is only a task-agnostic transformation of the coordinates $\mathbf{x}$, but the location embedding carries task-specific information. For example, it can contain information about spatial distributions of species if trained on geo-aware species fine-grained recognition tasks (mac2019presence, ; mai2023sphere2vec, ; mai2023csp, ; cole2023spatial, ; wu2024torchspatial, ; cole2023spatial, ; yang2022dynamic, ; liu2025gair, ). Some prior work, e.g., NeRF (mildenhall2020nerf, ), utilized positional encoding to represent location information. This task-agnostic method focuses on capturing the position or order of elements within a sequence. In contrast, many location encoders are specifically designed to capture context-aware location information (mai2020space2vec, ). Please refer to Appendix A.2 and mai2023sphere2vec for more details.
@@ 3 Preliminaries
Real Basis of Spherical Harmonics
Let $\mathbf{p}=(\theta,\phi)$ be a location on the spherical surface using angular coordinates where $\theta\in[0,\pi)$ and $\phi\in[0,2\pi)$. For any function $F(\theta,\phi)$ on the sphere, there exists a unique infinite-dimensional real-valued vector of coefficients $\{C_{lm}\}$ (we may call it coefficient vector) such that $\forall(\theta,\phi),F(\theta,\phi)=\sum_{l=0}^{\infty}\sum_{m=-l}^{l}C_{lm}Y_{lm}(\theta,\phi)$ where $l$ is called degree and $m$ is called order and $Y_{lm}(\theta,\phi)$ is the real basis of spherical harmonics at degree $l$ and order $m$. The detailed computation of $Y_{lm}$ can be found in Appendix A.3. In this way, any function on the sphere can be uniquely represented by its coefficient vector.
Spherical Dirac Delta Function
Similar to the Dirac delta function in Euclidian space we can define a spherical Dirac delta function is a probability density function over the spherical surface whose mass all concentrates on one point: $\delta_{(\theta_{0},\phi_{0})}(\theta,\phi)=\begin{cases}\infty&\theta=\theta_{0},\phi=\phi_{0}\\ 0&\text{otherwise}\end{cases}$. It can uniquely represent any point $(\theta_{i},\phi_{i})$ on the sphere by mapping it to $\delta_{(\theta_{i},\phi_{i})}$. Representing a point as a function allows us to use spherical harmonics to represent points on the spherical surface.
4 LocDiff Framework
In this section, we will introduce the theory and techniques we employ in our LocDiff model that enable spherical location generation via latent diffusion. Our aim is to find a position encoding space that does not suffer from the sparsity problem and the non-linearity problem so we can efficiently perform latent diffusion. We first analyze what properties we need to achieve this and propose the Spherical Harmonics Dirac Delta (SHDD) Encoding-Decoding framework accordingly. Then we prove that SHDD satisfies all the desired properties. Following that, we propose the Conditional Siren-UNet (CS-UNet) architecture to learn the conditional backward process for latent diffusion. We also develop computational techniques based on the properties of SHDD representation so that the training and inference of LocDiff are efficient.